[figure: 9.7 cm]

GROUND-STATE ENERGY OF A FRUSTRATED ± J ISING

MODEL ON A STRIP OF SQUARES

M.Błaszyk and B.Fechner Institute of Physics, A. Mickiewicz University, Poznań, Poland

We investigated a ±J Ising model on an infinite strip of squares of the width of one, two and three bonds with a random distribution of ferro- and antiferromagnetic exchange interactions [1-4]. The exact ground-state energy as function of the concentration of antiferromagnetic bonds was derived and its singular points in the complex plane were found.We calculated also the ground-state energy using two different procedures either of which can be applied to a strip of any width.

1) B. Derrida, J. Vannimenus and Y. Pomeau: J.Phys.C 11 (1978) 4749

2) T. Morita and T. Horiguchi: J.Phys.C 13 (1980) 6483

3) B. Fechner and M. Błaszyk: J.Phys.C 19 (1986) 2785

4) T. Kadowaki, Y. Nonomura and H. Nishimori: J.Phys.Soc.Jpn. 65 (1996) 1609

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M.Błaszyk

Address for correspondence : Wydział Fizyki UAM ul. Umultowska 85 61-614 Poznań

Email address : bfechner@main.amu.edu.pl